OpenStudy (anonymous):

HELP!: What is the equation of the line that passes through the points (2,1) and (4,5) in the Cartesian coordinate plane?

OpenStudy (anonymous):

\[y-y_{1}=\frac{y_{2}-y_{1}}{x_{2}-x{1}}(x-x_{1})\] x and y are variables of the equation of line, x1,x2,y1,y2 are given in question

OpenStudy (anonymous):

general equation of a line in rectangular coordinate system is y=mx+c m = slope c= line intersects y axis that point 2,1 5,4 y2-y1/x2-x1 = 1-4/2-5 = -3/-3 = 1 y=x+c m=1 1=2+c c=-1 y=x-1 ans

OpenStudy (anonymous):

A. y = 23 x − 43 B. y = 23 x − 13 C. y = 23 x + 73 D. y = 23 x + 4